(A) Find the slope of the line that passes through the given points. (B) Find the point-slope form of the equation of the line. (C) Find the slope-intercept form of the equation of the line. (D) Find the standard form of the equation of the line. (5,6) and (5,12) (A) Choose the correct answer for the slope below. OA, m=□ (Type an integer or a simplified fraction.) O B. The slope is not defined.
A plant can manufacture 50 golf clubs per day at a total daily cost of $5403 and 60 golf clubs per day for a total cost of $6153 (A) Assuming that daily cost and production are linearly related, find the total daily cost, C, of producing x golf clubs. (B) Graph the total daily cost for 0 s xs 200. (C) Interpret the slope and y intercept of the cost equation. (A) C (Do not include the $ symbol in your answer.)
A farmer buys a new tractor for $155,000 and assumes that it will have a trade-in value of $78,000 after 10 years. The farmer uses a constant rate of depreciation to determine the annual value of the tractor (A) Find a linear model for the depreciated value V of the tractor t years after it was purchased.
We found that the marketing research department for the company that manufactures and sells memony chips for microcomputers established the following price-demand and revenue functions: p(x) 105-4x R(x) xp(x) x(105-4x) where p(x) is the wholesale price in dollars at which x million chips can be sold, and R(x) is in millions of dollars. Both functions have domain 1sxs21. Price-demand function Revenue function (A) Which of the following is the graph of the revenue function? O C. 2005 10 15 20 400 400 200- 0 10 15 20 20-15-10-5 0 Similar Question Click to select your answer and then click Check Answer. parts remaining Final Checlk Clear All
At a price of $2.25 per bushel, the supply of a certain grain is 7100 million bushels and the demand is 7800 million bushels. At a price of $2.34 per bushel, the supply is 7500 million bushels and the demand is 7700 million bushels. (A) Find a price-supply equation of the form p mx+ b, where p is the price in dollars and x is the supply in millions of bushels (B) Find a price-demand equation of the form p mx+ b, where p is the price in dollars and x is the demand in millions of bushels. (C) Find the equilibrium point. (D) Graph the price-supply equation, price-demand equation, and equilibrium point in the same coordinate system (A) The price-supply equation is p (Type an exact answer) Enter your answer in the answer box and then click Check Answer. parts Clear All Check Answer remaining
Answer #1

Solution:

\small \left ( A \right ) Option  \small B. The slope is not defined

\small \left ( x_{1} ,y_{1}\right )\equiv \left ( 5,6 \right ),\: \: \: \: \left ( x_{2},y_{2} \right )\equiv \left ( 5,12 \right )

\small m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}=\frac{12-6}{5-5}=\frac{6}{0}

\small \left ( B \right )

The point -slope form is

\small y-y_{1}=m\left ( x-x_{1} \right )

\small y-6=\frac{12}{0}\left ( x-5 \right )

\small \therefore x-5=0

\small \therefore x=5

\small \left ( C \right )

The slope-intercept of the equation of the line is  \small x=5

\small \left ( D \right )

The standard form is  \small 0\cdot y+x=5